An astronaut leaves earth and travels at 0.92c. He spends 15 years exploring space. When he returns, how much time will have passed on earth ?
1
Expert's answer
2019-04-08T09:45:08-0400
"t'=\frac{t}{\sqrt{1-(\frac{v}{c})^2}}"

"t'=\frac{15}{\sqrt{1-0.92^2}}=38 years"